A ceiling fan rotates at an angular speed of 30.0 revolution per minute. Find
(a) the angular speed of the fan in rad/s.
(b) the speed of the tip of the fan blade which is 0.60 m from the axis of rotation.
(c) the angular displacement of a blade of the fan in 5.0 s.
1
Expert's answer
2020-10-30T13:12:23-0400

a) The angular speed is: "f = 30rpm = 0.5Hz". Then the angular speed in rad/s is:


"\omega = 2\pi f = 2\pi\cdot 0.5 = \pi\space rad/s"

b) The linear speed of points "r = 0.6m" from the axis of rotation is given by:


"v = \omega r = \pi\cdot 0.6 = 0.6\pi\space m/s"


c) The angular displacement of a blade is given by:


"\varphi = \omega t = \pi\space rad/s \cdot 5s = 5\pi \space rad"

Answer. a) "\pi\space rad/s", b) "0.6\pi\space m/s", c) "5\pi \space rad".